OpenStudy (anonymous):

HELP!!! Evaluate in cylindrical coordinates. Q is bounded by z=x^2+y^2 and z=8. Triple integral below....

OpenStudy (anonymous):

\[\int\limits_{Q}^{Q}\int\limits_{Q}^{Q}\int\limits_{Q}^{Q}10(x ^{2}+y ^{2})\sqrt{x ^{2}+y ^{2}}dV\]

ganeshie8 (ganeshie8):

where are u stuck ?

OpenStudy (anonymous):

I'm not sure how to start the integral without the bounds